Part I

Elementary Probabilities

[170/99%]
Show
LaTeX
Problem 1

Let Ω\Omega denote the totality of students at the University of Athens and A1,A2,A3,A4A_{1}, A_{2}, A_{3}, A_{4}, the sets of freshmen, sophomores, juniors, and seniors, respectively. Moreover, let FF denote the set of female students and CC the set of Cypriot students. Express in words each of the following sets:

?
(a)

(A1∪A2)∩F(A_{1} \cup A_{2}) \cap F;

(b)

FC′FC';

(c)

A1F′CA_{1}F'C;

(d)

A3FC′A_{3}FC';

(e)

(A1∪A2)CF(A_{1} \cup A_{2})CF.

Problem 2

Give the simplified forms of the sets:

?
(a)

(A∪B)∩(A∪C)(A \cup B) \cap (A \cup C);

(b)

(A∪B)∩(A′∪B)(A \cup B) \cap (A' \cup B);

(c)

(A∪B)∩(A′∪B)∩(A′∪B′)(A \cup B) \cap (A' \cup B) \cap (A' \cup B').

Problem 3

Express each of the following events in terms of the events AA, BB, and CC, and the operations of complementation, union, and intersection:

?
(a)

at least one of the events AA, BB, CC occurs;

(b)

at most one of the events AA, BB, CC occurs;

(c)

none of the events AA, BB, CC occurs;

(d)

all three events occur;

(e)

exactly one of the events AA, BB, CC occurs;

(f)

AA and BB occur but not CC;

(g)

AA occurs, if not then BB does not occur either.

Give the Venn diagram for each of the above.

Problem 4

Let Ω\Omega denote the sample space corresponding to the chance experiment of tossing a coin three times. Let AA be the event that heads appear exactly twice, let BB be the event that at least two heads appear, and let CC the event that heads appear when tails have appeared at least once.

?
(a)

Give the elements of AA, BB, CC;

(b)

Describe the events: (i) A′BA'B; (ii) A′B′A'B'; (iii) ACAC.

Problem 5

In each of the exercises 5--8, give the number of points in the sample space Ω\Omega and the points of the events defined therein.

A family has 4 children. Let the events: AA: boys and girls alternate, BB: the first and fourth child are boys, CC: as many boys as girls, and DD: three successive children of the same sex.

?
Problem 6

A salesman is arranging his schedule for visiting each of three towns aa, bb, cc, twice. AA: the first and last visit are in aa.

?
Problem 7

Certain diseases cause cell disorders which can be distinguished into four categories. Blood tests are conducted on each of 4 patients and in each case the category is noted. Let AA: the patients belong to the same category, BB: two patients belong to the same category.

?
Problem 8

An elevator carries two persons and stops at three floors. Let AA: they get off at different floors, BB: one gets off at the first floor.

?
Problem 9

In how many ways can a lady having 10 dresses, 5 pairs of shoes, and 2 hats be dressed?

?
Problem 10

In how many ways can we place in a bookcase two works each of three volumes and two works each of four volumes, so that the volumes of the same work are not separated?

?
Problem 11

In how many ways can rr objects be distributed to nn persons if there is no restriction on the number of objects that a person may receive?

?
Problem 12

In how many ways can 5 boys and 5 girls be seated around a table so that no 2 boys sit next to each other?

?
Problem 13

Eight points are chosen on the circumference of a circle. How many chords can be drawn by joining these points in all possible ways? If the 8 points are considered vertices of a polygon, how many triangles and how many hexagons can be formed?

?
Problem 14

In how many ways can nn persons be seated around a table if 2 arrangements are regarded as the same when each person has the same right and left neighbors? At a dinner at which the nn persons are seated randomly, what is the probability that a specific husband is seated next to his wife?

?
Problem 15

In how many ways can 20 recruits be distributed into 4 groups each consisting of 5 recruits? In how many ways can they be distributed into 4 camps, each camp receiving 5 recruits?

?
Problem 16

Using 7 consonants and 5 vowels, how many words consisting of 4 consonants and 3 vowels can we form?

?
Problem 17

Show that for every xx and for every positive integer nn:

  1. (xn−1)+(xn)=(x+1n)\binom {x}{n-1} + \binom {x}{n} = \binom {x+1}{n};

  2. (xn)=(x−1n−1)+(x−2n−1)+⋯+(n−1n−1)\binom {x}{n} = \binom {x-1}{n-1} + \binom {x-2}{n-1} + \cdots + \binom {n-1}{n-1}.

?
Problem 18

If nn, mm, rr, are positive integers prove that

(m0)(n−mr)+(m1)(n−mr−1)+⋯+(mr)(n−m0)=(nr). \binom {m}{0}\binom {n-m}{r} + \binom {m}{1}\binom {n-m}{r-1} + \cdots + \binom {m}{r}\binom {n-m}{0} = \binom {n}{r}.
?
Problem 19

For every integer n≥2n \geq 2, show the following:

  1. 1−(n1)+(n2)+⋯+(−1)n(nn)=01 - \binom {n}{1} + \binom {n}{2} + \cdots + (-1)^{n}\binom {n}{n} = 0;

  2. (n1)+2(n2)+3(n3)+⋯=n2n−1\binom {n}{1} + 2\binom {n}{2} + 3\binom {n}{3} + \cdots = n2^{n-1};

  3. 2⋅1⋅(n2)+3⋅2⋅(n3)+4⋅3⋅(n4)+⋯=n(n−1)2n−22 \cdot 1 \cdot \binom {n}{2} + 3 \cdot 2 \cdot \binom {n}{3} + 4 \cdot 3 \cdot \binom {n}{4} + \cdots = n(n-1)2^{n-2}.

?
Problem 20

Using Pascal's triangle, show that for every aa and positive integers rr, nn:

  1. ∑k=0n(a−kr)=(a+1r+1)−(a−nr+1)\displaystyle \sum_{k=0}^{n} \binom {a-k}{r} = \binom {a+1}{r+1} - \binom {a-n}{r+1};

  2. ∑k=0n(−1)k(ak)=(−1)n(a−1n)\displaystyle \sum_{k=0}^{n} (-1)^{k}\binom {a}{k} = (-1)^{n}\binom {a-1}{n}.

?
Problem 21

Using the result of Exercise 18, show that

(n0)2+(n1)2+⋯+(nn)2=(2nn). \binom {n}{0}^{2} + \binom {n}{1}^{2} + \cdots + \binom {n}{n}^{2} = \binom {2n}{n}.
?
Problem 22

Using the result of Exercise 21, prove that

∑ν=0n(2n)!(ν!)2[(n−ν)!]2=(2nn)2. \sum _{\nu =0}^{n} \frac{(2n)!}{(\nu !)^{2}[(n-\nu )!]^{2}} = \binom {2n}{n}^{2}.
?
Problem 23

Using the binomial theorem, show by induction that

∑k=1n(nk)(−1)k−1k=∑k=1n1k. \sum _{k=1}^{n} \binom {n}{k}\frac{(-1)^{k-1}}{k} = \sum _{k=1}^{n} \frac{1}{k}.
?
Problem 24

As in Exercise 23, show that for positive integers rr and nn

∑k=0n(k+r−1r−1)=(n+rr). \sum _{k=0}^{n} \binom {k+r-1}{r-1} = \binom {n+r}{r}.

Show that this is a special case of Problem 20(ii).

?
Problem 25

In the expansion of (1+x)n(1+x)^{n} where x>0x > 0 and nn is a positive integer, let aka_{k} be the term containing xkx^{k}. Find the values of kk for which aka_{k} becomes maximum. What is the maximum term of (1+e)100(1+e)^{100}?

?
Problem 26

Given P(A)=1/3\mathbb {P}\left(A\right) = 1/3, P(B)=1/4\mathbb {P}\left(B\right) = 1/4, P(AB)=1/6\mathbb {P}\left(AB\right) = 1/6, find the following probabilities:

P(A′),P(A′∪B),P(A∪B′),P(AB′),P(A′∪B′). \mathbb {P}\left(A'\right), \quad \mathbb {P}\left(A' \cup B\right), \quad \mathbb {P}\left(A \cup B'\right), \quad \mathbb {P}\left(AB'\right), \quad \mathbb {P}\left(A' \cup B'\right).
?
Problem 27

Given P(A)=3/4\mathbb {P}\left(A\right) = 3/4 and P(B)=3/8\mathbb {P}\left(B\right) = 3/8, show that:

?
(a)

P(A∪B)≥3/4\mathbb {P}\left(A \cup B\right) \geq 3/4,

(b)

1/8≤P(AB)≤3/81/8 \leq \mathbb {P}\left(AB\right) \leq 3/8. Give inequalities analogous to (a) and (b) for P(A)=1/3\mathbb {P}\left(A\right) = 1/3 and P(B)=1/4\mathbb {P}\left(B\right) = 1/4.

Problem 28

For any two events AA and BB show that:

P(AB)−P(A)P(B)=P(A′)P(B)−P(A′B)=P(A)P(B′)−P(AB′). \mathbb {P}\left(AB\right) - \mathbb {P}\left(A\right)\mathbb {P}\left(B\right) = \mathbb {P}\left(A'\right)\mathbb {P}\left(B\right) - \mathbb {P}\left(A'B\right) = \mathbb {P}\left(A\right)\mathbb {P}\left(B'\right) - \mathbb {P}\left(AB'\right).
?
Problem 29

By induction show that for arbitrary events A1,A2,…,AnA_{1}, A_{2}, \ldots , A_{n}, the following inequality (due to Bonferoni) holds:

P(A1A2⋯An)≥∑i=1nP(Ai)−(n−1)=1−∑i=1nP(Aic). \mathbb {P}\left(A_{1}A_{2} \cdots A_{n}\right) \geq \sum _{i=1}^{n} \mathbb {P}\left(A_{i}\right) - (n-1) = 1 - \sum _{i=1}^{n} \mathbb {P}\left(A_{i}^{c}\right).
?
Problem 30

For any three events AA, BB, and CC, show that (cf. formula (1.1), Poincaré's theorem for nn events A1,…,AnA_{1}, \ldots , A_{n}: P(A1∪A2∪⋯∪An)=S1−S2+⋯+(−1)n−1Sn\mathbb {P}\left(A_{1} \cup A_{2} \cup \cdots \cup A_{n}\right) = S_{1} - S_{2} + \cdots + (-1)^{n-1}S_{n}, where for each k=1,2,…,nk = 1, 2, \ldots , n, Sk=∑i1<i2<⋯<ikP(Ai1Ai2⋯Aik)S_{k} = \sum_{i_{1} < i_{2} < \cdots < i_{k}} \mathbb {P}\left(A_{i_{1}}A_{i_{2}} \cdots A_{i_{k}}\right), the summation being over all combinations of the nn events taken kk at a time)

P(A∪B∪C)=P(A)+P(B)+P(C)−P(AB)−P(AC)−P(BC)+P(ABC). \begin{align} \mathbb {P}\left(A \cup B \cup C\right) & = \mathbb {P}\left(A\right) + \mathbb {P}\left(B\right) + \mathbb {P}\left(C\right) - \mathbb {P}\left(AB\right) - \mathbb {P}\left(AC\right) \\ & \quad - \mathbb {P}\left(BC\right) + \mathbb {P}\left(ABC\right). \end{align}

Application: The percentages of students who passed courses AA, BB, and CC are as follows: AA: 50, BB: 40, CC: 30, AA and BB: 35, BB and CC: 20, and 15 passed all three courses. What is the percentage of students who succeeded in at least one of the three courses?

?
Problem 31

A box contains balls numbered 1,2,…,n1, 2, \ldots , n. A ball is drawn at random:

?
(a)

What is the probability that its label number is divisible by 3 or 4?

(b)

Examine the case in (a) as n→∞n \rightarrow \infty.

Problem 32

In terms of P(A)\mathbb {P}\left(A\right), P(B)\mathbb {P}\left(B\right), P(C)\mathbb {P}\left(C\right), P(AB)\mathbb {P}\left(AB\right), P(AC)\mathbb {P}\left(AC\right), P(BC)\mathbb {P}\left(BC\right), and P(ABC)\mathbb {P}\left(ABC\right) express, for k=0,1,2,3k = 0, 1, 2, 3, the probabilities that:

  1. exactly kk of the events AA, BB, and CC, occur;

  2. at least kk of the events AA, BB, CC occur.

?
Problem 33

By induction prove Poincaré's theorem, i.e., formula (1.1).

?
Problem 34

If d(A,B)=P(A△B)d(A,B) = \mathbb {P}\left(A \triangle B\right), show that dd has all the properties of a distance function; the symmetric difference A△BA \triangle B between two sets AA and BB is defined by A△B=AB′∪A′BA \triangle B = AB' \cup A'B.

?
Problem 35

Three winning tickets are drawn from an urn of 100 tickets. What is the probability of winning for a person who buys: (a) 4 tickets? (b) only one ticket?

?
Problem 36

A bakery makes 80 loaves of bread daily. Ten of them are underweight. An inspector weighs 5 loaves at random. What is the probability that an underweight loaf will be discovered?

?
Problem 37

Find the probability that among seven persons:

?
(a)

no two were born on the same day of the week (Sunday, Monday, etc.);

(b)

at least two were born on the same day;

(c)

two were born on a Sunday and two on a Tuesday.

Problem 38

A group of 2N2N boys and 2N2N girls is randomly divided into two equal groups. What is the probability that each group has the same number of boys and girls?

?
Problem 39

The coefficients aa, bb, cc of the quadratic equation ax2+bx+c=0ax^{2} + bx + c = 0 are determined by throwing a die three times. Find the probabilities that: (a) the roots are real; (b) the roots are complex.

?
Problem 40

In the game of poker a ``hand of cards'' means 5 cards randomly selected (without replacement) from a deck of 52 cards. What is the probability that a hand of cards:

?
(a)

consists of an ace, a queen, a jack, a king, and a ten of the same suit?

(b)

contains 4 cards of the same denomination (aces, etc.)?

(c)

consists of cards with consecutive values, except aces, jacks, kings, and queens?

Problem 41

In bridge, a hand of cards'' consists of 13 cards drawn at random (without replacement) from a deck of 52 cards (the 52 cards are equally distributed to four players). Find the probabilities that a hand of cards'' will contain:

  1. v1v_{1} clubs, v2v_{2} spades, and v3v_{3} diamonds;

  2. vv aces (v=0,1,…,4)(v = 0, 1, \ldots , 4);

  3. v1v_{1} aces and v2v_{2} kings. In a bridge game find the probabilities that:

?
(a)

each player has an ace;

(b)

some player has all the aces;

(c)

some player has v1v_{1} aces and his partner v2v_{2} aces (v1+v2≤4)(v_{1} + v_{2} \leq 4).

Problem 42

Six girls are to enter a dance with 10 boys to form a ring so that every girl is between two boys:

?
(a)

What is the probability that some specified boy remains between 2 boys?

(b)

A spectator notices that a certain girl enters next to a certain boy. Is it random?

Problem 43

Five letters are selected at random one after another from the 26 letters of the English alphabet; (a) with replacement, (b) without replacement. Find, for each of the cases (a) and (b), the probabilities that the word formed: (i) contains an a'', (ii) consists of vowels, (iii) is the word woman''.

?
(a)
(b)
Problem 44

Prove that the number of ways in which rr indistinguishable balls can be distributed in nn cells (or, equivalently, the number of different solutions of the equation x1+x2+⋯+xn=rx_{1} + x_{2} + \cdots + x_{n} = r where xi≥0x_{i} \geq 0 (i=1,2,…,n)(i = 1, 2, \ldots , n) are integers), equals

(n+r−1n−1)=(n+r−1r). \binom {n+r-1}{n-1} = \binom {n+r-1}{r}.
?
Problem 45

Consider rr indistinguishable balls randomly distributed in nn cells (Bose--Einstein statistics). What is the probability that exactly mm cells remain empty?

?
Problem 46

From an ordinary deck of 52 cards, cards are drawn successively until an ace appears. What is the probability that the first ace will appear: (a) at the nnth draw? (b) after the nnth card?

?
(a)
(b)
Problem 47

In a classroom there are vv students.

?
(a)

What is the probability that at least two students have the same birthday?

(b)

What is the minimum value of vv which secures probability 1/21/2 that at least two have a common birthday?

Problem 48

NN letters are placed at random in NN envelopes. Show that the probability that each letter will be placed in a wrong envelope is ∑k=2N(−1)k(1k!)\displaystyle \sum_{k=2}^{N} (-1)^{k} \left(\frac{1}{k!}\right).

?
Problem 49

An urn contains nrnr balls numbered 1,2,…,n1, 2, \ldots , n in such a way that rr balls bear the same number ii for each i=1,2,…,ni = 1, 2, \ldots , n. NN balls are drawn at random without replacement. Find the probability that,

?
(a)

exactly mm of the numbers will appear in the sample,

(b)

each of the nn numbers will appear at least once.

Problem 50

(Continuation of Problem 49.) Balls are drawn until each of the numbers 1,2,…,n1, 2, \ldots , n appears at least once. What is the probability that mm balls will be needed?

?
Problem 51

NN men run out of a men's club after a fire and each takes a coat and a hat. Prove that:

?
(a)

the probability that no one will take his own coat and hat is

∑k=1N(−1)k(N−k)!N! k!; \sum _{k=1}^{N} (-1)^{k} \frac{(N-k)!}{N! \, k!};
(b)

the probability that each man takes a wrong coat and a wrong hat is

[∑k=2N(−1)k1k!]2. \left[\sum _{k=2}^{N} (-1)^{k} \frac{1}{k!}\right]^{2}.
Problem 52

Suppose every packet of the detergent TIDE contains a coupon bearing one of the letters of the word TIDE. A customer who has all the letters of the word gets a free packet. All the letters have the same possibility of appearing in a packet. Find the probability that a housewife who buys 8 packets will get:

  1. one free packet,

  2. two free packets.

?
Problem 53

An urn contains N1N_{1} white and N2N_{2} black balls. When two balls are randomly drawn the probability that both be white is 1/21/2.

?
(a)

What is the minimum value of N1N_{1}?

(b)

What is the minimum value of N1N_{1} when N2N_{2} is an even number?

(c)

What is the minimum value of the total number N=N1+N2N = N_{1} + N_{2} of balls in the urn?

Problem 54

If the events AA, BB are independent, show that the pairs (A,B′)(A, B'), (A′,B)(A', B), (A′,B′)(A', B') also consist of independent events.

?
Problem 55

If the events AA, BB, and CC are mutually independent, then the pairs (A,BC)(A, BC), (B,AC)(B, AC), (C,AB)(C, AB) also consist of independent events.

?
Problem 56

Suppose that for the independent events AA, BB, and CC we have P(A)=a\mathbb {P}\left(A\right) = a, P(A∪B∪C)=1−b\mathbb {P}\left(A \cup B \cup C\right) = 1 - b, P(ABC)=1−c\mathbb {P}\left(ABC\right) = 1 - c, and P(A′B′C)=x\mathbb {P}\left(A'B'C\right) = x. Prove that the probability xx satisfies the equation

ax2+[ab−(1−a)(a−c−1)]x+b(1−a)(1−c)=0. ax^{2} + \left[ab - (1 - a)(a - c - 1)\right]x + b(1 - a)(1 - c) = 0.

Hence conclude that

c>(1−a)2+ab1−a. c > \frac{(1 - a)^{2} + ab}{1 - a}.

Moreover, show that

P(B)=(1−c)(x+b)ax,P(C)=xx+b. \mathbb {P}\left(B\right) = \frac{(1 - c)(x + b)}{ax}, \qquad \mathbb {P}\left(C\right) = \frac{x}{x + b}.
?
Problem 57

Let the events A1,A2,…,AnA_{1}, A_{2}, \ldots , A_{n} be independent and P(Ai)=p\mathbb {P}\left(A_{i}\right) = p (i=1,2,…,ni = 1, 2, \ldots , n). What is the probability that:

?
(a)

at least one of the events will occur?

(b)

at least mm of the events will occur?

(c)

exactly mm of the events will occur?

Problem 58

A die is thrown as long as necessary for an ace or a 6 to turn up. Given that no ace turned up at the first two throws, what is the probability that at least three throws will be necessary?

?
Problem 59

A parent particle can be divided into 0, 1, or 2 particles with probabilities 1/41/4, 1/21/2, 1/41/4, respectively. It disappears after splitting. Beginning with one particle, the progenitor, let us denote by XiX_{i} the number of particles in the ii generation. Find,

?
(a)

P(X2>0)\mathbb {P}\left(X_{2} > 0\right),

(b)

the probability that X1=2X_{1} = 2 given that X2=1X_{2} = 1.

Problem 60

An urn contains nn balls numbered 1,2,…,n1, 2, \ldots , n. We select at random rr balls,

?
(a)

with replacement,

(b)

without replacement.

What is the probability that the largest selected number is mm?

Problem 61

Two athletic teams AA and BB play a series of independent games until one of them wins 4 games. The probability of each team winning in each game equals 1/21/2. Find the probability that the series will end,

?
(a)

in at most 6 games,

(b)

in 6 games given that team AA won the first two games.

Problem 62

It is suspected that a patient has one of the diseases A1A_{1}, A2A_{2}, A3A_{3}. Suppose that the population percentages suffering from these illnesses are in the ratio 2:1:12:1:1. The patient is given a test which turns out to be positive in 25%25\% of the cases of A1A_{1}, 50%50\% of A2A_{2}, and 90%90\% of A3A_{3}. Given that out of three tests taken by the patient two were positive, find the probability for each of the three illnesses.

?
Problem 63

The population of Nicosia (Cyprus) is 75%75\% Greek and 25%25\% Turkish. 20%20\% of the Greeks and 10%10\% of the Turks speak English. A visitor to the town meets someone who speaks English. What is the probability that he is a Greek? Interpret your answer in terms of the population of the town.

?
Problem 64

Two absent-minded room mates, mathematicians, forget their umbrellas in some way or another. AA always takes his umbrella when he goes out, while BB forgets to take his umbrella with probability 1/21/2. Each of them forgets his umbrella at a shop with probability 1/41/4. After visiting three shops they return home. Find the probability that:

?
(a)

they have both umbrellas;

(b)

they have only one umbrella;

(c)

BB has lost his umbrella given that there is only one umbrella after their return.

Problem 65

Consider families of nn children and let AA be the event that a family has children of both sexes, and let BB be the event that there is at most one girl in the family. Show that the only value of nn for which the events AA and BB are independent is n=3n = 3, assuming that each child has probability 1/21/2 of being a boy.

?
Problem 66

At the college entrance examination each candidate is admitted or rejected according to whether he has passed or failed the test. Of the candidates who are really capable, 80%80\% pass the test; and of the incapable, 25%25\% pass the test. Given that 40%40\% of the candidates are really capable, find the proportion of capable college students.

?
Problem 67

AA and BB throw alternately a pair of dice in that order. AA wins if he scores 6 points before BB gets 7 points, in which case BB wins. If AA starts the game what is his probability of winning?

?
Problem 68

Three players P1P_{1}, P2P_{2}, and P3P_{3} throw a die in that order and, by the rules of the game, the first one to obtain an ace will be the winner. Find their probabilities of winning.

?
Problem 69

NN players A1,A2,…,ANA_{1}, A_{2}, \ldots , A_{N} throw a biased coin whose probability of heads equals pp. A1A_{1} starts (the game), A2A_{2} plays second, etc. The first one to throw heads wins. Find the probability that AkA_{k} (k=1,2,…,Nk = 1, 2, \ldots , N) will be the winner.

?
Problem 70

Urn AA contains w1w_{1} white balls and b1b_{1} black balls. Urn BB contains w2w_{2} white balls and b2b_{2} black balls. A ball is drawn from AA and is placed into BB, and then a ball is transferred from BB to AA. Finally, a ball is selected from AA. What is the probability that the ball will be white?

?
Problem 71

Ten percent of a certain population suffer from a serious disease. A person suspected of the disease is given two independent tests. Each test makes a correct diagnosis 90%90\% of the time. Find the probability that the person really has the illness:

?
(a)

given that both tests are positive;

(b)

given that only one test is positive.

Problem 72

A player randomly chooses one of the coins AA, BB. Coin AA has probability of heads 3/43/4, and coin BB has probability of heads 1/41/4. He tosses the coin twice.

?
(a)

Find the probability that he obtains: (i) two heads, (ii) heads only once.

(b)

Instead of the above procedure, suppose the player can choose an unbiased (symmetric) coin, which he tosses twice. Which procedure should he follow in order to maximize the probability of at least one head?

Problem 73

Urn AA contains 5 black balls and 6 white balls, and urn BB contains 8 black balls and 4 white balls. Two balls are transferred from BB to AA and then a ball is drawn from AA.

?
(a)

What is the probability that this ball is white?

(b)

Given that the ball drawn is white, what is the probability that at least one white ball was transferred to AA?

Problem 74

The Pap test makes a correct diagnosis with probability 95%95\%. Given that the test is positive for a lady, what is the probability that she really has the disease? How do you interpret this? Assume that one in every 2,000 women, on average, has the disease.

?
Problem 75

A secretary goes to work following one of three routes AA, BB, CC. Her choice of route is independent of the weather. If it rains, the probabilities of arriving late, following AA, BB, CC, are 0.060.06, 0.150.15, 0.120.12, respectively. The corresponding probabilities, if it does not rain, are 0.050.05, 0.100.10, 0.150.15.

?
(a)

Given that on a sunny day she arrives late, what is the probability that she took route CC? Assume that, on average, one in every four days is rainy.

(b)

Given that on a day she arrives late, what is the probability that it is a rainy day?

Problem 76

At an art exhibition there are 12 paintings of which 10 are original. A visitor selects a painting at random and before he decides to buy, he asks the opinion of an expert about the authenticity of the painting. The expert is right in 9 out of 10 cases on average.

?
(a)

Given that the expert decides that the painting is authentic, what is the probability that this is really the case?

(b)

If the expert decides that the painting is a copy, then the visitor returns it and chooses another one; what is the probability that his second choice is an original?

Problem 77

What is the conditional probability that a hand at poker (see Exercise 40) consists of spades, given that it consists of black cards?

Note: Of the 52 cards of an ordinary deck, the 13 spades and 13 clubs are black whereas the 13 hearts and 13 diamonds are red.

?
Problem 78

The probability pkp_{k} that a family has kk children is given by p0=p1=ap_{0} = p_{1} = a, pk=(1−2a)2−(k−1)p_{k} = (1 - 2a)2^{-(k - 1)} (k≥2k \geq 2). It is known that a family has two boys. What is the probability that:

?
(a)

the family has only two children?

(b)

the family has two girls as well?

Problem 79

An amplifier may burn one or both of its fuses if one or both of its tubes are defective. Let the events:

Ai:only tube i is defective (i=1,2); A_{i}: \text{only tube } i \text{ is defective } (i = 1, 2); A3:both tubes are defective; A_{3}: \text{both tubes are defective}; Bj:fuse j burns out (j=1,2); B_{j}: \text{fuse } j \text{ burns out } (j = 1, 2); B3:both fuses burn out. B_{3}: \text{both fuses burn out}.

The conditional probabilities of burning the fuses (given the state of the tubes, i.e., the P(Bj∣Ai)\mathbb {P}\left( B_{j} \mid A_{i} \right)), appear in the following table:

i/jB1B2B3A10.70.20.1A20.30.60.1A30.20.20.6 \begin{array}{c|ccc} i / j & B_{1} & B_{2} & B_{3} \\ \hline A_{1} & 0.7 & 0.2 & 0.1 \\ A_{2} & 0.3 & 0.6 & 0.1 \\ A_{3} & 0.2 & 0.2 & 0.6 \end{array}

Find the probabilities that:

?
(a)

both tubes are defective given that both fuses were burnt out;

(b)

only tube 2 is defective given that both fuses were burnt out;

(c)

only tube 2 is defective given that at least one fuse was burnt out.

The events A1A_{1}, A2A_{2}, A3A_{3} have probabilities 0.30.3, 0.20.2, 0.10.1, respectively.

Problem 80

An urn contains 7 white balls numbered 1,2,…,71, 2, \ldots , 7 and 3 black balls numbered 8,9,108, 9, 10. Five balls are randomly selected,

?
(a)

with replacement;

(b)

without replacement.

For each of the cases (a) and (b) give the distribution:

  1. of the number of white balls in the sample;

  2. of the minimum number in the sample;

  3. of the maximum number in the sample;

  4. of the minimum number of balls needed for selecting a white ball.

Problem 81

(Continuation of Problem 80.) An urn contains 7 white balls numbered 1,2,…,71, 2, \ldots , 7 and 3 black balls, now numbered 1,21, 2, and 33 (instead of 8,9,108, 9, 10 as in Problem 80). Five balls are randomly selected, (a) with replacement, (b) without replacement. As in Problem 80, give for each of the cases (a) and (b) the distribution:

  1. of the number of white balls in the sample;

  2. of the minimum number in the sample;

  3. of the maximum number in the sample;

  4. of the minimum number of balls needed for selecting a white ball.

?
Problem 82

A machine normally makes items of which 4%4\% are defective. Every hour the producer draws a sample of size 10 for inspection. If the sample contains no defective items he does not stop the machine. What is the probability that the machine will not be stopped when it has started producing items of which 10%10\% are defective.

?
Problem 83

One per thousand of a population is subject to certain kinds of accident each year. Given that an insurance company has insured 5,000 persons from the population, find the probability that at most 2 persons will incur this accident.

?
Problem 84

A certain airline company, having observed that 5%5\% of the persons making reservations on a flight do not show up for the flight, sells 100 seats on a plane that has 95 seats. What is the probability that there will be a seat available for every person who shows up for the flight?

?
Problem 85

Workers in a factory incur accidents at the rate of two accidents per week. Calculate the probability that there will be at most two accidents,

  1. during 1 week,

  2. during 2 weeks,

  3. in each of 2 weeks.

?
Problem 86

Suppose that the suicide rate in a certain state is four suicides per one million inhabitants per month. Find the probability that in a certain town of population 500,000 there will be at most four suicides in a month. Would you find it surprising that during 1 year there were at least 2 months in which more than four suicides occurred?

?
Problem 87

A coin is tossed 30 times. For n=1,2,…,20n = 1, 2, \ldots , 20, what is the conditional probability that exactly 10+n10+n heads appear given that the first 10 tosses resulted in heads. Show that the conditional probability that 10+n10+n tosses will result in heads, given that heads appeared at least ten times, equals

(3010+n)12n∑k=020(3010+k)12k. \frac{\displaystyle \binom {30}{10+n}\frac{1}{2^{n}}}{\displaystyle \sum _{k=0}^{20} \binom {30}{10+k}\frac{1}{2^{k}}}.
?
Problem 88

How many children should a family have so that with probability 0.95 it has at least a boy and at least a girl.

?
Problem 89

Show that the Poisson probabilities p(k∣λ)p(k \mid \lambda ) satisfy the recurrence relation

p(k∣λ)=λkp(k−1∣λ), p(k \mid \lambda ) = \frac{\lambda }{k}p(k-1 \mid \lambda ),

and hence determine the values of kk for which the terms p(k∣λ)p(k \mid \lambda ) reach their maximum (for given λ\lambda).

?
Problem 90

Verify that each of the following functions ff is a probability density function and sketch its graph.

?
(a)

f(x)=1−∣1−x∣for 0<x<2\displaystyle f(x) = 1-\left|1-x\right| \qquad \text{for } 0 < x < 2

(b)

f(x)=1πββ2+(x−a)2for −∞<x<∞,\displaystyle f(x) = \frac{1}{\pi }\frac{\beta }{\beta^{2}+(x-a)^{2}} \qquad \text{for } -\infty < x < \infty ,

(c)

f(x)=12σe−(∣x−μ∣)/σfor −∞<x<∞,\displaystyle f(x) = \frac{1}{2\sigma }e^{-(\left|x-\mu \right|)/\sigma } \qquad \text{for } -\infty < x < \infty ,

(d)

f(x)=14xe−x/2for 0<x<∞.\displaystyle f(x) = \tfrac {1}{4}xe^{-x/2} \qquad \text{for } 0 < x < \infty .

Problem 91

The amount of bread (in hundreds of kilos) that a bakery sells in a day is a random variable with density

f(x)={cxfor 0≤x<3,c(6−x)for 3≤x<6,0otherwise. f(x) = \begin{cases} cx & \text{for } 0 \leq x < 3, \\ c(6-x) & \text{for } 3 \leq x < 6, \\ 0 & \text{otherwise.} \end{cases}
  1. Find the value of cc which makes ff a probability density function.

  2. What is the probability that the number of kilos of bread that will be sold in a day is, (a) more than 300 kilos? (b) between 150 and 450 kilos?

  3. Denote by AA and BB the events in (a) and (b), respectively. Are AA and BB independent events?

?
Problem 92

Suppose that the duration in minutes of long-distance telephone conversations follows an exponential density function;

f(x)=15e−x/5for x>0. f(x) = \tfrac {1}{5}e^{-x/5} \qquad \text{for } x > 0.

Find the probability that the duration of a conversation:

?
(a)

will exceed 5 minutes;

(b)

will be between 5 and 6 minutes;

(c)

will be less than 3 minutes;

(d)

will be less than 6 minutes given that it was greater than 3 minutes.

Problem 93

A number is randomly chosen from the interval (0,1)(0, 1). What is the probability that:

?
(a)

its first decimal digit will be a 1;

(b)

its second decimal digit will be a 5;

(c)

the first decimal digit of its square root will be a 3?

Problem 94

The height of men is normally distributed with mean μ=167\mu = 167 cm and standard deviation σ=3\sigma = 3 cm.

  1. What is the percentage of the population of men that have height,

    1. greater than 167 cm,

    2. greater than 170 cm,

    3. between 161 cm and 173 cm?

  2. In a random sample of four men what is the probability that:

    1. all will have height greater than 170 cm;

    2. two will have height smaller than the mean (and two bigger than the mean)?

?
Problem 95

A machine produces bolts the length of which (in centimeters) obeys a normal probability law with mean 5 and standard deviation σ=0.2\sigma = 0.2. A bolt is called defective if its length falls outside the interval (4.8,5.2)(4.8, 5.2).

?
(a)

What is the proportion of defective bolts that this machine produces?

(b)

What is the probability that among ten bolts none will be defective?

Problem 96

Consider a shop at which customers arrive at random at a rate of twenty persons per hour. What is the probability that the time intervals between succesive arrivals will be:

?
(a)

shorter than 3 minutes;

(b)

longer than 4 minutes.

(c)

Suppose that 10%10\% of the customers buy a certain object. Find the distribution of the number of customers who buy an object in an hour.

Problem 97

If XX is a continuous random variable with cumulative distribution function FF and density function ff, show that the random variable Y=X2Y = X^{2} is also continuous and express its cumulative distribution function and density in terms of FF and ff.

?
Problem 98

(Continuation of Problem 97.) Find the density of Y=X2Y = X^{2} when XX has:

?
(a)

the normal distribution N(μ,σ2)N(\mu , \sigma^{2});

(b)

the Laplace distribution (see (2.9));

(c)

the Cauchy distribution (see (2.12)).

Problem 99

As in Exercises 97 and 98 with Y=∣X∣Y = \left|X\right|.

?
(a)
(b)
(c)
Problem 100

If the log⁡X\log X is normally distributed then XX is said to have a lognormal distribution. Find its density.

?
Problem 101

Show that the integral ∫−∞+∞∣x−m∣f(x) dx\int_{-\infty }^{+\infty } \left|x - m\right| f(x)\, dx becomes minimum when mm is the median of the distribution with density ff.

?
Problem 102

Let f(x)f(x) denote the density function of the random variable XX. Suppose that XX has a symmetric distribution about aa, that is, f(x+a)=f(a−x)f(x + a) = f(a - x) for every xx. Show that the mean E[X]\mathbb {E}\left[X\right] equals aa, provided it exists.

?
Problem 103

Show that E[XY]2≤E[X2]E[Y2]\mathbb {E}\left[XY\right]^{2} \leq \mathbb {E}\left[X^{2}\right] \mathbb {E}\left[Y^{2}\right] provided the second moments exist.

?
Problem 104

Show that ∣E[X]∣≤E[∣X∣]\left|\mathbb {E}\left[X\right]\right| \leq \mathbb {E}\left[\left|X\right|\right].

?
Problem 105

For every set AA of real numbers, we define the indicator function by

IA(x)={1,if x∈A,0,if x∉A. I_{A}(x) = \begin{cases} 1, & \text{if } x \in A, \\ 0, & \text{if } x \notin A. \end{cases}

Show that

P(A)=E[IA(X)]=∫AdFX(x). \mathbb {P}\left(A\right) = \mathbb {E}\left[I_{A}(X)\right] = \int _{A} dF_{X}(x).
?
Problem 106

If E[X]=E[X2]=0\mathbb {E}\left[X\right] = \mathbb {E}\left[X^{2}\right] = 0, show that P(X=0)=1\mathbb {P}\left(X = 0\right) = 1.

?
Problem 107

Show that the mean μ\mu of a random variable XX has the property

min⁡cE[(X−c)2]=E[(X−μ)2]=Var⁡[X]. \min _{c} \mathbb {E}\left[(X - c)^{2}\right] = \mathbb {E}\left[(X - \mu )^{2}\right] = \operatorname {Var}\left[X\right].
?
Problem 108

Show that for a continuous random variable XX with density function ff and cumulative distribution function FF

μ=E[X]=∫0+∞[1−F(x)] dx−∫−∞0F(x) dx. \mu = \mathbb {E}\left[X\right] = \int _{0}^{+\infty } [1 - F(x)]\, dx - \int _{-\infty }^{0} F(x)\, dx.

Consequently, in the graph below μ=area⁡(A)−area⁡(B)\mu = \operatorname {area}(A) - \operatorname {area}(B).

?
Problem 109

If the nnth moment of the random variable XX with distribution function FF exists, show that

E[(X−c)k]=k∫c∞(x−c)k−1[1−F(x)] dx−k∫−∞c(x−c)k−1F(x) dx,1≤k≤n. \mathbb {E}\left[(X - c)^{k}\right] = k \int _{c}^{\infty } (x - c)^{k-1} [1 - F(x)]\, dx - k \int _{-\infty }^{c} (x - c)^{k-1} F(x)\, dx, \qquad 1 \leq k \leq n.
?
Problem 110

Show that the first nn moments determine the first nn central moments and, conversely, that the first nn central moments with the mean determine the first nn moments.

?
Problem 111

If XX is bounded, i.e., there is a constant k<∞k < \infty such that P(∣X∣≤k)=1\mathbb {P}\left(\left|X\right| \leq k\right) = 1, then XX has moments of every order.

?
Problem 112

If the nn-order moment μn′\mu_{n}' exists, then show that there exist all μk′\mu_{k}' (k=1,2,…,n−1k = 1, 2, \ldots , n - 1).

?
Problem 113

Show that a necessary condition for the mean of a random variable with distribution function FF to exist is that

lim⁡x→−∞xF(x)=lim⁡x→+∞x[1−F(x)]=0. \lim _{x \to -\infty } x F(x) = \lim _{x \to +\infty } x [1 - F(x)] = 0.
?
Problem 114

In a lottery, nn numbers are selected from the NN numbers 1,2,…,N1, 2, \ldots , N. Find the variance of the sum SnS_{n} of the selected numbers.

?
Problem 115

An urn contains N1N_{1} white balls and N2N_{2} black balls; nn balls are drawn at random,

?
(a)

with replacement,

(b)

without replacement.

What is the expected number of white balls in the sample?

Problem 116

A student takes a multiple-choice test consisting of two problems. The first one has 3 possible answers and the second one has 5. The student chooses, at random, one answer as the right one from each of the two problems. Find:

?
(a)

the expected number, E[X]\mathbb {E}\left[X\right], of the right answers XX of the student;

(b)

the Var⁡[X]\operatorname {Var}\left[X\right].

Generalize.

Problem 117

In a lottery that sells 3,000 tickets the first lot wins $1,000, the second $500, and five other lots that come next win $100 each. What is the expected gain of a man who pays 1 dollar to buy a ticket?

?
Problem 118

A die is thrown until the result ``ace or even number'' appears three times. Find the expected number of throws:

?
(a)

in one performance of the experiment;

(b)

in ten repetitions.

Problem 119

AA pays 1 dollar for each participation in the following game: three dice are thrown; if one ace appears he gets 1 dollar, if two aces appear he gets 2 dollars and if three aces appear he gets 8 dollars; otherwise he gets nothing. Is the game fair, i.e., is the expected gain of the player zero? If not, how much should the player receive when three aces appear to make the game fair?

?
Problem 120

A player has 15 dollars. If heads appear on the first toss he receives 1 dollar and he withdraws from the game. If tails appear he bets 2 dollars and if he wins the second toss he withdraws; otherwise he continues playing on a bet of 4 dollars. If he loses, he bets the remaining amount of 8 dollars. What is the expected gain of the player?

?
Problem 121

Of the parcels mailed abroad 10 never reach their destination. Two books may be sent separately or in a single parcel. Each book is worth $2. The postage for each book sent separately is 10¢ and for both books in a single parcel 15¢. For each of the two ways of mailing, find:

?
(a)

the probability that both books reach their destination;

(b)

the probability that at least one book reaches its destination;

(c)

the expected net (after postage) value of the commodities reaching their destination.

For each of the three criteria (a), (b), and (c), which way of mailing is preferable?

Problem 122
?
(a)

Two indentical coins are thrown once in such a way that if one shows heads so does the other (dependent). Let P(heads)=p\mathbb {P}\left(\text{heads}\right) = p and XX be the total number of heads that appear. Find the mean and the variance of the random variable XX.

(b)

Two different coins with probabilities of heads p1p_{1} and p2p_{2} are thrown independently. Find the mean and the variance of the random variable XX as defined in (a).

Problem 123

Let XX be a random variable with distribution function

F(x)={1−0.8e−x,x≥0,0,x<0. F(x) = \begin{cases} 1 - 0.8 e^{-x}, & x \geq 0, \\ 0, & x < 0. \end{cases}

Plot the graph of F(x)F(x) and then evaluate E[X]\mathbb {E}\left[X\right]. Give an example of a random variable XX that has the above distribution.

?
Problem 124

If the random variable XX is N(μ,1)\mathscr {N}\left(\mu , 1\right), show that the random variable Y=[1−Φ(X)]/φ(X)Y = [1 - \Phi (X)]/\varphi (X), where Φ\Phi and φ\varphi denote the distribution function and the density of N(0,1)\mathscr {N}\left(0, 1\right), respectively, has mean value 1/μ1/\mu.

?
Problem 125

The truncated Poisson distribution with the zero class missing has probability function

P(X=k)=λk(eλ−1)k!,k=1,2,…. \mathbb {P}\left(X = k\right) = \frac{\lambda ^{k}}{(e^{\lambda } - 1) k!}, \qquad k = 1, 2, \ldots .

Find E[X]\mathbb {E}\left[X\right] and Var⁡[X]\operatorname {Var}\left[X\right].

?
Problem 126

Three players AA, BB, and CC play a game as follows: At the first stage AA and BB play against each other while CC stays out of the game. The winner plays against CC at the second stage. Then the winner of the second game plays against the last loser, and so on. The winner is declared the one who wins two games in succession. Find the elements (simple events) of the sample space Ω\Omega of all possible outcomes of the game. If the probability of each player winning a game is 1/21/2, determine the probabilities of the corresponding possible outcomes and show that their sum equals one.

?
Problem 127

A symmetrical coin is tossed until the same result appears twice in succession. Describe the sample space in terms of the results H (heads'') and T (tails''). Find the probabilities of the events:

?
(a)

the experiment terminates earlier or at the seventh trial;

(b)

an even number of trials is required.

What is the expected number of trials?

Problem 128

A pair of dice is tossed six times. What is the probability that all the faces will appear twice?

?
Problem 129

A deck of 52 cards is divided among four players so that each of them gets 13 cards (game of bridge). Find the probability that at least one player will have a complete suit (spades, hearts, diamonds, clubs).

?
Problem 130

Two coins C1C_{1} and C2C_{2} have a probability of falling heads p1p_{1} and p2p_{2}, respectively. You win a bet if in three tosses you get at least two heads in succession. You toss the coins alternately starting with either coin. If p1>p2p_{1} > p_{2}, what coin would you select to start the game?

?
Problem 131

Ten pairs of shoes are in a closet. Four shoes are selected at random. Find the probability that there will be at least one pair among the four shoes selected.

?
Problem 132

Let NN cells be numbered 1,2,…,N1, 2, \ldots , N. We randomly throw balls into them. The process is continued until a ball falls in the cell bearing the number 1. What is the probability that:

?
(a)

nn throws will be necessary?

(b)

more than nn throws will be necessary?

Problem 133

(Continuation of Problem 132.) The process terminates when any one of the cells receives two balls.

?
(a)

What is the probability that more than nn throws will be necessary?

(b)

What is the expected number of throws?

Problem 134

From a usual deck of cards we draw one card after another. What is the probability that:

?
(a)

the nnth card will be the first ace?

(b)

the first ace will appear among the first nn cards?

Problem 135

The 52 cards of an ordinary deck of cards are placed successively one after the other and from left to right. Find the probability that the thirteenth spade will appear before the thirteenth diamond.

?
Problem 136

Three balls are drawn at random one after another and without replacement from nn balls numbered 1,2,…,n1, 2, \ldots , n. Find the probability that the first will bear a number smaller than that of the second ball.

?
Problem 137

An urn contains mm kinds of objects all in the same proportion. Objects are drawn with replacement, one after another, until each kind appears at least once. What is the probability pνp_{\nu } that ν\nu objects will be required?

?
Problem 138

A newspaper dealer gets nn papers every day for sale. The number XX of papers sold is a random variable following the Poisson distribution with parameter λ\lambda. For each paper sold he earns 1 cent, for each unsold paper he loses bb cents. Let YY denote the net gain of the dealer. Find:

?
(a)

E[Y]\mathbb {E}\left[Y\right];

(b)

How many newspapers should he get every day in order to maximize his profit?

Problem 139

Three numbers are selected at random one after another and without replacement from nn numbers 1,2,…,n1, 2, \ldots , n. What is the probability that the first number drawn will be the smallest and the second number the largest?

?
Problem 140

The events AA, BB, and CC are independent with P(A)=0.2\mathbb {P}\left(A\right) = 0.2, P(B)=0.3\mathbb {P}\left(B\right) = 0.3, and P(C)=0.1\mathbb {P}\left(C\right) = 0.1. Find the probability that at least two will occur among the three events.

?
Problem 141

We throw four coins simultaneously and we repeat it once more. What is the probability that the second throw will result in the same configuration as the first throw when the coins are:

?
(a)

distinguishable;

(b)

indistinguishable.

Problem 142

An urn contains ww white balls and bb black balls. We draw balls successively and without replacement one after another until a white ball appears for the first time. Let XX be the required number of draws. Find:

?
(a)

the distribution of XX;

(b)

E[X]\mathbb {E}\left[X\right].

(c)

Using the result of (a) show the identity

1+n−wn−1+(n−w)(n−w−1)(n−1)(n−2)+⋯+(n−w)⋯2⋅1(n−1)⋯(w+1)w=nw, 1+\frac{n-w}{n-1}+\frac{(n-w)(n-w-1)}{(n-1)(n-2)}+\cdots +\frac{(n-w) \cdots 2 \cdot 1}{(n-1) \cdots (w+1) w} = \frac{n}{w},

where n=w+bn = w+b.

Problem 143

(Continuation of Problem 142.) Let w=b=Nw = b = N and draw balls successively and without replacement one after another until all the 2N2 N balls are drawn. What is the probability that in a stage of the experiment the same number of white and black balls has been drawn?

?
Problem 144

Let B1,…,Bν,…,B_{1}, \ldots , B_{\nu }, \ldots , be a partition of the sample space Ω\Omega and P(C)>0\mathbb {P}\left(C\right) > 0. Show the following theorem of total probability:

P(A∣C)=∑j=1∞P(Bj∣C)P(A∣BjC). \mathbb {P}\left( A \mid C \right) = \sum _{j=1}^{\infty } \mathbb {P}\left( B_{j} \mid C \right) \mathbb {P}\left( A \mid B_{j} C \right).
?
Problem 145

Each of the three Misses T. C---Rena, Nike, and Galatea---wants to accompany me on a trip. Since I cannot take all of them, I play the following game: I say to them, ``I will think of one of the numbers 1, 2, 3. Rena, you guess first. If you guess right you will come with me; otherwise Nike will guess next. If she guesses right she will come with me. Otherwise Galatea will accompany me''. Galatea complains that the game is unfair. Is she right? Suppose every one of them writes down on a piece of paper the number which she guesses I thought of and I follow the above sequence checking Rena's number first, etc. This is repeated until one is chosen to come with me. Galatea and Nike protest this time. Are they justified?

?
Problem 146

A die is thrown ν\nu times. Find the probability that each of kk (1≤k≤61 \leq k \leq 6) given faces of the die appears at least once.

?
Problem 147

The folded normal distribution has density

f(x)=c12πe−x2/2,x>0. f(x) = c \frac{1}{\sqrt{2 \pi }} e^{-x^{2} / 2}, \qquad x > 0.

Determine the constant cc and then evaluate the mean of the distribution.

?
Problem 148

Suppose the lifetime of an electric lamp of a certain type obeys a normal law with mean μ=180\mu = 180 hours and standard deviation σ=20\sigma = 20. In a random sample of four lamps:

?
(a)

What is the probability that all four lamps have a lifetime greater than 200 hours?

(b)

The random sample of four lamps is placed in an urn and then we randomly draw one lamp from it; what is the probability that the lamp will have a lifetime greater than 200 hours?

Generalize the conclusion.

Problem 149

The game of “craps” is played as follows. The gambler throws two dice. If at the first throw he gets 7 or 11 he wins, and if he gets 2, 3, or 12 he loses. For each of the other sums the game is continued in two ways:

?
(a)

the gambler continues throwing the two dice until he wins with a 7 or he loses with the results of the outcome of the first throw;

(b)

the gambler continues until he loses with 7 or wins with the result of the first throw.

What is the probability of the gambler winning in cases (a) and (b)?

Problem 150

Two gamblers AA and BB agree to play as follows. They throw two dice and if the sum SS of the outcomes is <10< 10, BB receives SS dollars from AA, otherwise BB pays AA xx dollars. Determine xx so that the game is fair.

?
Problem 151

An urn contains nn white balls and nn red balls. We draw two balls at first, and then another two, and so on, until all the balls are drawn. Find the probability that each of the selected pairs consists of a white ball and a red ball.

?
Problem 152

Each of two urns AA and BB contains nn balls numbered 1 to nn. We draw one ball from each of the urns. Find the probability that the ball drawn from AA bears a number smaller than that drawn from BB.

?
Problem 153

The rumor mongers (see Feller, 1957, p. 55). In a town of N+1N+1 inhabitants, a person tells a rumor to a second person, who in turn repeats it to a third person, and so on. At each step the recipient of the rumor is chosen at random from the NN inhabitants available.

  1. Find the probability that the rumor will be told nn times without, (a) returning to the originator, (b) being repeated to any person. Do the same problem when at each step the rumor is told to kk persons.

  2. In a large town where the rumor mongers constitute 100p%100 p \% of the population, what is the probability that the rumor does not return to the originator?

?
Problem 154

Pascal’s problem. In the game of tossing a fair coin, the first one to obtain nn successes (heads or tails) wins. Show that the game is fair (i.e., each gambler has a probability of winning equal to 1/21/2). Suppose that the game was interrupted when the first gambler had won kk tosses and the second gambler had won mm tosses (0≤k,m<n0 \leq k, m < n). Calculate for each gambler the probability of winning if the game is continued. Hence deduce how the stake should be divided after interrupting the game.

?
Problem 155

Chebyshev’s problem. Find the probability that a given fraction m/nm/n (where mm and nn are integers) is irreducible.

?
Problem 156

Let φ(n)\varphi (n) denote Euler's function, i.e., the number of (positive) integers which are primes relative to nn and smaller than nn. Using a probabilistic argument, show that

φ(n)=n∏p/n(1−1p), \varphi (n) = n \prod _{p / n} \left(1-\frac{1}{p}\right),

where the product extends over all prime divisors pp of nn.

?
Problem 157

Assume that the number of insect colonies in a certain area follows the Poisson distribution with parameter λ\lambda, and that the number of insects in a colony has a logarithmic distribution with parameter pp. Show that the total number of insects in the area is a negative binomial with parameters q=1−pq = 1-p and −λ/log⁡(1−p)-\lambda / \log (1-p).

?
Problem 158

Electronic tubes come in packages, each containing NN tubes. Let pkp_{k} denote the probability that a package contains kk defective tubes (0≤k≤m0 \leq k \leq m). A sample of nn tubes is taken from a package and it is observed that r≤mr \leq m tubes are defective.

?
(a)

What is the probability that the selected package actually contains kk (≥r\geq r) defective tubes?

(b)

If according to customer demand a package is considered not acceptable whenever it contains d≥rd \geq r defective tubes, what is the probability that the package is not accepted?

Problem 159

The probability pnp_{n} that nn customers visit a supermarket in one day is pn=pnqp_{n} = p^{n} q, n=0,1,…n = 0, 1, \ldots. Two out of three customers, on average, buy a certain type of item. The probability that an item is defective is 1/41/4.

?
(a)

What is the probability that a customer buys a nondefective item?

(b)

Given that kk nondefective items were sold, show that the conditional probability ana_{n} that nn customers visited the shop is given by

an=(nk)pn−k(2−p)k+1/2n+1. a_{n} = \binom {n}{k} p^{n-k}(2-p)^{k+1} / 2^{n+1}.
Problem 160

A manufacturer sells an item for $1. If the weight of the item is less than W0W_{0}, it cannot be sold and it represents a complete loss. The weight WW of an item follows the normal distribution N(μ,1)\mathscr {N}\left(\mu , 1\right); the cost cc per item is given by c=α+βWc = \alpha +\beta W (α\alpha, β\beta positive constants). Determine the mean μ\mu so that the expected profit is maximized.

?
Problem 161

A clerk lives at AA and works at CC and he starts work at 9 a.m. The clerk always takes the train from AA to BB which is supposed to reach BB at 8:40 a.m. Buses from BB leave for CC every 15 minutes and the bus which leaves at 8:45 a.m. is supposed to arrive at 8:56 a.m.

The train on average experiences delays of 2 minutes and has a standard deviation of 4 minutes. The bus always leaves on time, but arrives on average after a 2-minute delay and a standard deviation of 3 minutes. What is the probability that the clerk arrives late? The clerk's employer drives to his office; he leaves at 8:45 a.m. and the driving time to the office has a mean value of 12 minutes and a standard deviation of 2 minutes. Find the probability that:

?
(a)

both, clerk and employer, arrive late;

(b)

the employer arrives earlier than the clerk.

Assume that the distributions involved are normal.

Problem 162

Quality control. In a continuous manufacturing process the percentage of defective articles is pp. To maintain the quality of the product at a certain standard, the articles are examined, one by one, until a sequence of fixed length rr appears with no defective article. Then total inspection terminates and only a certain fraction ff of the product is chosen at random for inspection until a defective article appears when the above process of 100%100\% inspection continues. Under the described sampling scheme find:

?
(a)

the probability of a defective sequence, i.e., that a defective article is not followed by rr successive good articles;

(b)

the expected number of articles in a defective sequence. Hence deduce the expected number, (i) of defective sequences, (ii) of inspected articles, after a defective article;

(c)

the expected proportion of the product subject to inspection;

(d)

the expected percentage ρ\rho of defective articles going out for sale provided that any discovered defective article is replaced by a good one;

(e)

for given values of ff and rr, what value p∗p^{*}, say, of pp maximizes ρ\rho.

Problem 163

The dilemma of the convict. Three convicts AA, BB, and CC appeal for parole and the board decides to free two of them. The convicts are informed of this decision but are not told who the two to be set free are. The guard, who is AA's friend, knows which convicts are going to be set free. AA sees that it would not be right to ask the guard about his own fate, but thinks that he might ask for the name of one of the other two who will be set free. He supposes that before he asks the probability of his being set free is 2/32/3, and when the guard answers that BB is pardoned, for example, the probability that he will be set free is lessened to 1/21/2, because either AA and BB will be set free or BB and CC. Having thought about this, AA is afraid to ask the guard. Is his fear justified?

?
Problem 164

Collection of coupons. In each box of a given product there is a coupon with a number from 1 to 6. If a housewife succeeds in getting the series 1--6 she receives a free box of the product. How many boxes must she buy, on average, before she gets a free one?

?
Problem 165

In a row of 19 seats in an amphitheatre 10 male students and 9 female students sit at random. Find the expected number of successive pairs of seats in which a male student and a female student sit.

?
Remark.
?

We used the fact that the mean of a sum of random variables is equal to the sum of the mean values of the corresponding random variables, no matter whether the variables are (statistically) independent or not.

Problem 166

AA, BB, and CC are about to play the following fatal game. Each one has a gun and will shoot at his target---AA, BB, and CC taking turns at each other in that order---and it is assumed that any one who is hit will not be shot at again. The shots, one after another, continue until one of the players remains unhit. What strategy should AA follow?

?
Problem 167

Samuel Pepys (who was about to place a bet) asked Newton which of the following events AA, BB, or CC is more probable: AA---at least one six when 6 dice are thrown; BB---at least two sixes when 12 dice are thrown; and CC---at least three sixes when 18 dice are thrown. Which answer (the correct one) did Newton give?

?
Problem 168

Birthday holidays. The Worker's Legal Code in Erehwon specifies as a holiday any day during which at least one worker in a certain factory has a birthday. All other days are working days. How many workers must the factory employ so that the number of working man-days is maximized during the year?

?
Problem 169

Bertrand’s paradox. A chord ABAB is randomly chosen in a circle of radius rr. What is the probability that the length of ABAB is less than rr?

?
Remark.
?

The variety of solutions exhibited is due to insufficient description of the corresponding random experiment which determines the choice of the chord. Thus in (a) we essentially choose at random a point MM in the circle and require the probability that MM lies in a smaller homocentric cycle. In (b) the points AA and BB are randomly chosen on the circumference, whereas in (c) we can imagine a cylinder rolling on a diameter of the circle (say KK′KK') and the equally likely events correspond to equal intervals on the diameter (irrespective of their position on the diameter).

Problem 170

The neophyte at the horseraces. At a horserace a neophyte better, who is about to bet a certain amount, wants to choose the best horse. It is assumed that no two horses are the same. The better looks at the horses as they pass by, one after another, and he can choose as the winner any one of the horses, but he cannot bet on a horse he has already let pass. Moreover, he can tell whether any horse passing by is better or worse compared with the preceding ones. Suppose nn horses take part in the race. How should he proceed to choose the horse to bet on in order to maximize his probability of winning? What proportion of horses should he wait to pass by before he makes his bet when nn is large? (It is assumed that horses parade in front of the neophyte in a random order.)

?